### v2

# FAST: Foreground‑aware Diffusion with Accelerated Sampling Trajectory for Segmentation‑oriented Anomaly Synthesis

Abstract
Industrial anomaly segmentation relies heavily on pixel-level annotations, yet real-world anomalies are often scarce, diverse, and costly to label. Segmentation-oriented industrial anomaly synthesis (SIAS) has emerged as a promising alternative; however, existing methods struggle to balance sampling efficiency and generation quality. Moreover, most approaches treat all spatial regions uniformly, overlooking the distinct statistical differences between anomaly and background areas. This uniform treatment hinders the synthesis of controllable, structure-specific anomalies tailored for segmentation tasks. In this paper, we propose FAST, a foreground-aware diffusion framework featuring two novel modules: the Anomaly-Informed Accelerated Sampling (AIAS) and the Foreground-Aware Reconstruction Module (FARM). AIAS is a training-free sampling algorithm specifically designed for segmentation-oriented industrial anomaly synthesis, which accelerates the reverse process through coarse-to-fine aggregation and enables the synthesis of state-of-the-art segmentation-oriented anomalies in as few as 10 steps. Meanwhile, FARM adaptively adjusts the anomaly-aware noise within the masked foreground regions at each sampling step, preserving localized anomaly signals throughout the denoising trajectory. Extensive experiments on multiple industrial benchmarks demonstrate that FAST consistently outperforms existing anomaly synthesis methods in downstream segmentation tasks. We release the code in https://github.com/Chhro123/fast-foreground-aware-anomaly-synthesis.
1 Introduction
Motivation. Industrial anomaly segmentation plays a vital role in modern manufacturing, aiming to localize abnormal regions at the pixel level. Unlike traditional anomaly detection, which typically performs binary classification at the image or region level, anomaly segmentation requires more fine-grained and precise localization of abnormal patterns. However, real-world anomalies are inherently scarce, diverse, and non-repeatable, making it difficult to collect data that fully captures the range of possible abnormal types. Moreover, acquiring high-quality pixel-level annotations is labor-intensive and costly, especially in industrial scenarios. To address these limitations, recent studies have increasingly explored the use of synthetic anomalies to expand the training data space and improve downstream performance.
Limitations. Despite recent advances, current anomaly synthesis methods face three fundamental limitations that hinder their effectiveness for segmentation tasks xu2025survey . (i) Lack of controllability. Most existing methods provide limited control over the structure, location, or extent of synthesized anomalies. This limitation is particularly evident in GAN-based approaches niu2020defect ; zhang2021defect ; du2022new . These methods typically adopt a one-shot generation paradigm, offering little flexibility in specifying where and how anomalies should appear. (ii) Neglect of segmentation-relevant properties. Training-free methods such as patch replacement or texture corruption li2021cutpaste ; zavrtanik2021draem may produce visible anomalies, but the synthesized patterns often lack the structural consistency and complexity of real-world industrial anomalies, which are critical for improving segmentation performance. (iii) Uniform treatment of spatial regions and inefficiency. Although recent diffusion-based methods he2024anomalycontrol ; hu2024anomalyxfusion ; shi2025few have mitigated the above issues, they still treat all spatial regions uniformly during both forward and reverse processes, without explicitly modeling the distinct statistical properties of anomaly regions zhang2024realnet ; yao2024glad . This absence of region-aware modeling prevents the model from preserving abnormal regions throughout the synthesis trajectory. Moreover, these models typically require hundreds to thousands of denoising steps ho2020denoising ; song2021denoising , resulting in a significant computational cost. While recent training-free methods liu2022pseudo aim to accelerate sampling, they fail to incorporate anomaly-aware cues, making them less effective for segmentation-oriented industrial anomaly synthesis (SIAS). These limitations motivate the need for SIAS models that support controllable anomaly synthesis, explicit modeling of anomaly regions, and efficient, task-aligned sampling strategies.
[…]
Contributions. In summary, our contributions are three-fold: (1) To mitigate the inefficiency and semantic misalignment of existing diffusion sampling, we introduce a training-free Anomaly-Informed Accelerated Sampling (AIAS) strategy that aggregates multiple denoising steps into a small number of coarse-to-fine analytical updates. (2) To address the lack of persistent anomaly-region representation, we propose a Foreground-Aware Reconstruction Module (FARM) that reconstructs pseudo-clean anomalies and reintegrates anomaly-aware noise at each step. (3) To support segmentation-oriented industrial anomaly synthesis, we design FAST, a controllable and efficient model. Extensive experiments on MVTec-AD and BTAD datasets demonstrate that it significantly outperforms existing methods in downstream segmentation tasks.
2 Related work
Industrial Anomaly Synthesis. Industrial anomaly synthesis aims to mitigate the scarcity of labeled abnormal samples in real-world inspection scenarios. Existing methods can be categorized into hand-crafted and DL-based approaches. Hand-crafted methods typically apply training-free manipulations to normal images, such as patch pasting pei2023self ; schluter2022natural or external texture blending zavrtanik2021draem ; yang2023memseg ; zhang2023destseg from sources like DTD cimpoi14describing , but they suffer from distributional deviation and limited realism. Deep learning-based methods alleviate these limitations by learning from real anomaly patterns. GAN-based methods duan2023few ; wen2022new can synthesize visually realistic anomalies but lack fine-grained controllability over anomaly shape and location. Diffusion-based methods duan2023few ; hu2024anomalydiffusion ; yang2024defect offer stronger generative capacity via large-scale pretrained models, yet treat all regions uniformly and lack explicit control over anomaly localization, which is essential for segmentation. To this end, we propose FAST, which integrates foreground-aware reconstruction and efficient, segmentation-oriented anomaly synthesis into a unified diffusion framework.
Acceleration of Discrete-Time Diffusion Models. Diffusion models can be categorized into continuous-time and discrete-time frameworks. Continuous formulations lu2022dpm ; lu2022dpm2 ; zheng2025diffusion adopt SDE/ODE-based parameterizations and leverage high-order solvers for efficient sampling. In contrast, standard DDPMs ho2020denoising model a discrete-time Markov chain with fixed variance schedules and require thousands of iterative denoising steps. While continuous-time solvers achieve notable speedups, they rely on continuously parameterized noise or score functions, which requires reformulating training objectives or interface in discrete-time models. Therefore, various acceleration techniques have been developed specifically for discrete-time diffusion. Some methods modify the generative process to reduce steps: DDGAN xiao2022tackling integrates GAN-based decoding, TLDM zheng2023truncated and ES-DDPM lyu2022accelerating truncate the forward process, and Blurring Diffusion Models hoogeboom2023blurring operate in the frequency domain. However, these methods require retraining and show limited generalization. In contrast, training-free approaches such as DDIM song2021denoising , PLMS liu2022pseudo , and GGDM watson2022learning accelerate sampling without model modification. Yet, they treat all spatial regions uniformly and lack task-specific guidance essential for SIAS. Recent work like CUT sun2024cut introduces external prompts for localized control for anomalies, but at the cost of multiple iterations per sampling step. In comparison, FAST proposes a novel training-free strategy that aggregates multiple denoising steps into coarse-to-fine segments while injecting mask-aware structural guidance, enabling efficient SIAS.
3 Methods
FAST for Anomaly Segmentation. The proposed FAST framework is built upon the Latent Diffusion Model (LDM) rombach2022high of $T$ steps. For notational simplicity, we denote the encoded latent of the original image as $x_{0}$, and its predicted reconstruction from the network as $\hat{x}_{0}$. We define $x_{t_{s}}$ as the noisy latent at timestep $t_{s}$, and $\hat{x}_{t_{s}}$ as the FARM-adjusted, anomaly-aware latent at the same step. Let $\mathcal{M}\in\{0,1\}^{H\times W}$ denote the binary anomaly mask, and $t_{s},t_{e}]$ represent a coarse-to-fine segment in AIAS, where $t_{e}<t_{s}$. Fig. 1 illustrates a single forward-reverse process at step $t_{s}$. In the forward phase, noise is added up to timestep $t_{s}$, yielding a noisy latent $x_{t_{s}}$. FARM ($F_{\phi}$ in Algorithm 1) then predicts a pseudo-clean anomaly latent $\hat{x}_{0}^{\text{an}}$, and adds noise to it up to timestep $t_{s}$ to obtain an anomaly-aware latent $\hat{x}_{t_{s}}$, which aims to match the observed $x_{t_{s}}$ in masked regions during training. In the corresponding reverse process, we divide the full denoising process into $S$ segments, each spanning $t_{s},t_{e}]$. Within each segment, AIAS approximates the posterior transition using: $q(x_{t_{e}}\mid x_{t_{s}},\hat{x}_{0})$. This formulation aggregates multiple DDPM steps into a single numerical update. FARM is also applied to refine $x_{t_{e}}$, ensuring the preservation of anomaly cues throughout the reverse process. More details can be seen in Algorithms 1 and 2.In addition, for the textual conditioning component of LDM, we follow the configuration of Anomaly Diffusion hu2024anomalydiffusion ; more implementation details can be found there.
[…]
Algorithm 1 FAST Training
1:repeat
2:  $x_{0}\sim q(x_{0})$, $\mathcal{M}$, and weights $\lambda_{1}$, $\lambda_{2}$
3:  $t_{s}\sim\text{Uniform}(\{1,\dots,T\})$, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
4:  $x_{t_{s}}=\sqrt{\bar{\alpha}_{t_{s}}}x_{0}+\sqrt{1-\bar{\alpha}_{t_{s}}}\epsilon$
5:  $\hat{x}_{t_{s}}=\sqrt{\bar{\alpha}_{t_{s}}}F_{\phi}(x_{t_{s}},\mathcal{M})+\sqrt{1-\bar{\alpha}_{t_{s}}}\epsilon$
6:  Take gradient descent step on:
$\nabla_{\theta}\left\|\epsilon-\epsilon_{\theta}(\hat{x}_{t_{s}},t_{s})\right\|^{2}$
$\nabla_{\phi}\left\|(\mathcal{M}\odot x_{0}-F_{\phi}(x_{t_{s}},t_{s},\mathcal{M}))\right\|^{2}$
7:until converged
Algorithm 2 FAST Sampling
(Details are shown in Supplementary Material A.4)
1:Initialize $x_{T}\sim\mathcal{N}(0,\mathbf{I})$
2:for each segment $t_{s},t_{e}]$ from $T\rightarrow 0$ do
3:  $\hat{\epsilon}=\epsilon_{\theta}(x_{t_{s}},t_{s})$
[…]
5:  AIAS:
$\mathbf{x}_{t_{e}}=F_{\phi}(q(x_{t_{e}}\mid x_{t_{s}},\hat{x}_{0}),t_{e},\mathcal{M}))$
6:end for
7:return $x_{0}$
3.1 Anomaly-Informed Accelerated Sampling
Standard DDPM allows us to directly compute the marginal distribution of $x_{t}$ given a clean sample $x_{0}$ and additive noise $\epsilon$. Therefore, the one-step posterior distribution of $x_{t-1}$ can be expressed as:
$q(x_{t-1}\mid x_{t},x_{0})=\mathcal{N}(A_{t}x_{0}+B_{t}x_{t},\,\sigma_{t}^{2}\mathbf{I}),$ (1)
where the coefficients are derived from the variance schedule as follows:
$A_{t}=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}},\quad B_{t}=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}},\quad\sigma_{t}^{2}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t},$
and $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$. All are closed-form coefficients derived from a predefined noise schedule. In practice, the true sample $x_{0}$ is not accessible during inference, and is typically replaced by a model prediction $\hat{x}_{0}$ obtained via denoising estimation. Equation 1 thus serves as the foundation for approximate posterior sampling, provided that $\hat{x}_{0}$ is a sufficiently accurate estimate of the ground truth $x_{0}$.
Theoretically, if we assume $\hat{x}_{0}=x_{0}$ holds exactly (i.e., the prediction perfectly matches the ground-truth image), then the entire reverse process becomes fully deterministic and analytically tractable, with the only source of stochasticity being the injected noise at each step. In this idealized setting, the reverse sampling trajectory is fully governed by closed-form probabilistic transitions. This forms the basis for Lemma. 1 (For brevity, the full proof is provided in the Supplementary Material A.1).
Lemma 1 (Linear–Gaussian closure).
Let $\{x_{k}\}_{k=0}^{K}\subset\mathbb{R}^{d}$ satisfy the recursion
$x_{k-1}\;=\;C_{k}\,x_{k}\;+\;d_{k}\;+\;\varepsilon_{k},\quad\varepsilon_{k}\sim\mathcal{N}(0,\Sigma_{k}),\quad\varepsilon_{k}\!\perp\!\!\{\!x_{k},\varepsilon_{k+1},\ldots\!\},$ (2)
where $C_{k}\!\in\!\mathbb{R}^{d\times d}$, $d_{k}\!\in\!\mathbb{R}^{d}$, and $\Sigma_{k}\!\in\!\mathbb{R}^{d\times d}$ are deterministic. Then, for every integer $m$ with $1\!\leq\!m\!\leq\!k$, $x_{k-m}$ is again an affine–Gaussian function of $x_{k}$:
$x_{k-m}\;=\;\underbrace{\Bigl(\prod_{i=0}^{m-1}C_{k-i}\Bigr)}_{=:C^{(m)}_{k}}x_{k}\;+\;\underbrace{\sum_{i=0}^{m-1}\Bigl(\prod_{j=1}^{i}C_{k-j}\Bigr)d_{k-i}}_{=:d^{(m)}_{k}}\;+\;\varepsilon^{(m)}_{k}$, (3)
where
$\varepsilon^{(m)}_{k}\sim\mathcal{N}\!\bigl(0,\Sigma^{(m)}_{k}\bigr),\qquad\Sigma^{(m)}_{k}=\sum_{i=0}^{m-1}\Bigl(\prod_{j=1}^{i}C_{k-j}\Bigr)\Sigma_{k-i}\Bigl(\prod_{j=1}^{i}C_{k-j}\Bigr)^{\!\top}.$ (4)
While the ideal condition $\hat{x}_{0}=x_{0}$ rarely holds in practice, the following properties justify the use of $\hat{x}_{0}$ in the multi-step formulation:
  • (i)
    The training objective of standard DDPM is explicitly designed to minimize the discrepancy between the predicted noise and the true noise. Consequently, the denoising model $\epsilon_{\theta}(x_{t},t)$ implicitly learns to reconstruct a close approximation of $x_{0}$ through the reverse reparameterization formula.
  • (ii)
[…]
Therefore, it is reasonable to treat $\hat{x}_{0}$ as fixed within a short temporal window. Under this assumption, multiple single-step reverse transitions can be analytically composed into a single multi-step affine–Gaussian kernel. This approximation and Lemma. 1 form the basis for Theorem 2, which characterizes the closed-form reverse process from $t_{s}$ to $t_{e}$ (For brevity, the full proof is provided in the Supplementary Material A.2 ).
Lemma 2 (Closed-form reverse from $\boldsymbol{t}_{s}\to t_{e}$}).
Fix indices $0\leq t_{e}<t_{s}\leq T$, and let the single-step coefficients $(A_{t},B_{t},\sigma_{t}^{2})$ be defined as in Eq. 13. Then the aggregated reverse kernel over $t_{s}\to\cdots\to t_{e}$ is affine–Gaussian:
$x_{t_{e}}=\Pi_{t_{e}}^{t_{s}}\,x_{t_{s}}+\Sigma_{t_{e}}^{t_{s}}\,\hat{x}_{0}+\varepsilon_{t_{e}},$ (5)
where
$\Pi_{t_{e}}^{t_{s}}:=\prod_{i=t_{e}+1}^{t_{s}}B_{i},\quad\Sigma_{t_{e}}^{t_{s}}:=\sum_{i=t_{e}+1}^{t_{s}}A_{i}\prod_{j=i+1}^{t_{s}}B_{j},\quad\varepsilon_{t_{e}}\sim\mathcal{N}\!\left(0,\sum_{i=t_{e}+1}^{t_{s}}\left(\prod_{j=i+1}^{t_{s}}B_{j}\right)^{2}\sigma_{i}^{2}\mathbf{I}\right).$
Therefore, it can be observed that in the limited segments (e.g., $t_{s}\to t_{e}$), there are the three scalars $\bigl(\Pi_{t_{e}}^{t_{s}},\Sigma_{t_{e}}^{t_{s}},\varepsilon_{t_{e}}\bigr)$, allowing us to precompute them once and re-use them during sampling. Lemma. 2 enables theoretical computation of posterior transitions between any two timesteps $t_{s}$ and $t_{e}$, allowing multi-step sampling in a manner distinct from DDIM. However, while the affine–Gaussian transition provides an efficient coarse approximation for the reverse path $x_{t_{s}}\to x_{t_{e}}$, the approximation may introduce residual artifacts in parctice. It is caused by the strong noise attenuation and the fixed $\hat{x}_{0}$ assumption. Moreover, since $x_{t}$ inherently entangles both the foreground and the background content, direct sampling through the affine-Gaussian kernel will ignore the critical spatial structure discrepancies for SIAS.
To better preserve anomaly-localized information while ensuring smooth global composition, we explicitly decompose the clean sample $x_{0}$ into two disjoint components:
$\mathbf{x}_{0}=\mathbf{x}_{0}^{\mathrm{an}}+\mathbf{x}_{0}^{\mathrm{bg}},\vskip-8.0pt$ (6)
where $\mathbf{x}_{0}^{\mathrm{an}}$ is the anomaly-only region (masked by $\mathcal{M}$), and $\mathbf{x}_{0}^{\mathrm{bg}}$ is the background. The background is independently forward-diffused:
$\mathbf{x}_{t_{e}}^{\mathrm{bg}}\sim q(\mathbf{x}_{t_{e}}^{\mathrm{bg}}\mid\mathbf{x}_{0}^{\mathrm{bg}}),$ (7)
while the anomaly foreground is refined by the learned FARM module (introduced later in Sec. 3.2), and merged with the background through spatial masking:
$\mathbf{x}_{t_{e}}^{R}$ $=\text{FARM}(\mathbf{x}_{t_{e}}),$ $\mathbf{x}_{t_{e}}$ $=\mathcal{M}\odot\mathbf{x}_{t_{e}}^{R}+(1-\mathcal{M})\odot\mathbf{x}_{t_{e}}^{\mathrm{bg}}.$ (8)
This foreground-aware fusion ensures consistent noise levels between anomalous and normal regions at each step, preserving local anomaly salience while maintaining global visual coherence. In practice, we also introduce a final fine-grained refinement stage using standard DDPM posterior sampling for small $t$ (e.g., $t=1$ or $t=2$) to restore the alignment between the coarse trajectory and the ground-truth posterior, and to enhance fine-scale texture fidelity. The complete sampling algorithm is summarized in Algorithm 3.
3.2 Foreground-Aware Reconstruction Module
Figure 2: The architecture of FARM. Given noisy latent $x_{t_{s}}$ and mask $\mathcal{M}$, the encoder $f_{\mathrm{enc}}$ extracts features $z_{t_{s}}$, which is also modulated by a background-adaptive soft mask $\tilde{\mathcal{M}}$ and related timestep embedding $\boldsymbol{\tau}_{t_{s}}$. The decoder $f_{\mathrm{dec}}$ then reconstructs the anomaly-only latent $\hat{x}_{0}^{\mathrm{an}}$, which is forward-diffused to produce anomaly-aware noise.
As discussed above, conventional diffusion models treat all spatial regions uniformly, which limits their ability to synthesize localized anomalies. To address this, we propose the Foreground-Aware Reconstruction Module (FARM), which reconstructs clean anomaly-only content from noisy latent inputs under both temporal and spatial guidance. As illustrated in Fig. 2, FARM adopts an encoder–decoder architecture. The encoder $f_{\mathrm{enc}}$ extracts deep representations from the noisy latent $x_{t_{s}}$, while the decoder $f_{\mathrm{dec}}$ progressively upsamples and integrates the binary mask $\mathcal{M}$ at multiple resolutions, ensuring spatial alignment with anomaly regions throughout the hierarchy.
To encode temporal context, we initialize sinusoidal timestep embeddings $\boldsymbol{\tau}_{t_{s}}\in\mathbb{R}^{d}$ and project them into latent space via a learned linear layer. These embeddings are added to the encoder output, modulating feature responses based on the current noise level and allowing the decoder to reconstruct temporally consistent structures.
In addition, to modulate background activation, we introduce a background-adaptive soft mask:
$\tilde{\mathcal{M}}=\mathcal{M}_{d}+(1-\mathcal{M}_{d})\cdot\sigma(f_{\mathrm{bg}}(\boldsymbol{\tau}_{t_{s}})),\vskip-10.0pt$ (9)
where $\mathcal{M}_{d}$ is a downsampled binary mask aligned with encoder resolution, and $f_{\mathrm{bg}}$ is a lightweight MLP. This design allows FARM to suppress irrelevant background features while adapting to the current timestep.
The encoded feature is computed as:
$z_{t_{s}}=\tilde{\mathcal{M}}\cdot f_{\mathrm{enc}}(\mathbf{x}_{t_{s}})+\text{Proj}(\boldsymbol{\tau}_{t_{s}}),$ (10)
and decoded into an anomaly-only latent: $\hat{x}_{0}^{\mathrm{an}}=f_{\mathrm{dec}}(z_{t_{s}},\mathcal{M})$.
To inject anomaly-aware noise into the sampling trajectory, the reconstructed anomaly is forward-diffused:
$\hat{x}_{t_{s}}^{\mathrm{an}}=\sqrt{\bar{\alpha}_{t_{s}}}\cdot\hat{x}_{0}^{\mathrm{an}}+\sqrt{1-\bar{\alpha}_{t_{s}}}\cdot\epsilon,\quad\epsilon\sim\mathcal{N}(0,\mathbf{I}),$ (11)
and replaces the original noise in masked regions:
$\hat{x}_{t_{s}}=(1-\mathcal{M})\cdot x_{t_{s}}+\mathcal{M}\cdot\hat{x}_{t_{s}}^{\mathrm{an}}.\vskip-10.0pt$ (12)
During training, FARM is supervised to ensure that the reconstructed anomalies match the masked regions of the noisy inputs. During inference, temporal and spatial guidance together enable FARM to introduce localized and temporally coherent anomaly signals into the reverse trajectory, ensuring alignment with the global generative process while enhancing fine-grained control.
4 Experiments
4.1 Implementation Details.
Datasets. We evaluate FAST on two widely-used industrial anomaly segmentation benchmarks: MVTec-AD bergmann2019mvtec and BTAD btad . For each anomaly class, we synthesize image–mask pairs using normal images, binary masks, and text prompts describing anomaly semantics. A total of 500 samples are generated per class, with approximately one-third used for training and the remainder reserved for evaluation. Evaluation Metrics. We report performance using mean Intersection over Union (mIoU) and pixel-wise accuracy (Acc), following standard practice in anomaly segmentation. Baselines. FAST is compared against six representative anomaly synthesis approaches: CutPaste li2021cutpaste , DRAEM zavrtanik2021draem , GLASS glass , the GAN-based SOTA method DFMGAN dfmgan , and diffusion-based SOTA models Anomaly Diffusion hu2024anomalydiffusion and RealNet zhang2024realnet . To simulate realistic deployment scenarios, we pair all generation methods with lightweight segmentation networks, including Segformer xie2021segformer , BiSeNet V2 yu2021bisenet , and STDC stdc . As our method adopts the same prompt-driven synthesis setup as AnomalyDiffusion hu2024anomalydiffusion , we omit the details here for brevity. Full specifications of the textual configuration, as well as other implementation details, including dataset preprocessing, sampling schedules, loss weights, and hyperparameter settings, are provided in the Supplementary Materials A.5.
4.2 Comparison Studies
Table 1: Evaluation of pixel-level segmentation accuracy on extended MVTec data using real-time Segformer. Detailed per-category results for other real-time segmentation model, such as BiseNet V2 and STDC are reported in Supplementary Material A.6.
Category CutPaste DRAEM GLASS DFMGAN RealNet AnomalyDiffusion FAST
mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$
bottle 75.11 79.49 79.51 84.99 70.26 76.30 75.45 80.39 77.96 83.90 76.39 83.54 86.86 90.90
cable 55.40 60.49 64.52 70.77 58.81 62.32 62.10 64.87 62.51 69.27 62.49 74.48 73.71 77.94
capsule 35.15 40.29 51.39 62.32 34.12 38.04 41.29 15.83 46.76 51.91 37.73 44.72 63.22 71.12
carpet 66.34 77.59 72.57 81.28 70.11 77.56 71.33 83.69 68.84 79.15 64.67 73.59 73.84 83.53
grid 29.90 46.72 47.75 67.85 37.43 46.30 37.73 54.13 37.55 48.86 38.70 51.82 52.45 70.70
hazel_nut 56.95 60.72 84.22 89.74 55.51 57.43 83.43 86.03 60.18 63.49 59.33 67.48 90.81 94.79
leather 57.23 63.49 64.12 71.49 62.05 73.38 60.96 68.02 68.29 77.16 56.45 62.51 66.60 74.18
metal_nut 88.78 90.94 93.51 96.10 88.15 90.52 92.77 94.93 91.28 94.09 88.00 91.10 94.65 96.88
pill 43.28 47.11 46.99 49.76 41.52 43.54 87.19 90.05 47.32 58.31 83.21 89.00 90.17 94.07
screw 25.10 31.35 46.96 59.03 35.94 42.37 46.65 50.79 47.12 55.17 38.47 49.49 49.94 57.48
tile 85.33 91.60 89.21 93.74 85.67 90.28 88.87 91.96 83.53 87.30 84.29 89.72 90.13 93.77
toothbrush 39.40 63.93 65.35 79.43 53.75 60.46 61.00 70.50 57.68 72.03 48.68 64.41 74.98 88.63
transistor 65.03 71.05 59.96 62.18 29.28 30.67 73.56 78.48 63.71 66.79 79.27 91.74 91.80 94.50
wood 49.64 60.47 67.52 73.28 50.91 53.16 67.00 80.84 61.84 89.54 60.16 74.62 78.77 86.31
zipper 65.39 71.89 69.29 79.36 69.98 79.31 66.34 70.50 68.78 78.50 65.36 72.66 72.80 84.73
Average 55.87 63.81 66.86 74.75 56.23 61.44 67.71 72.07 62.89 71.70 62.88 72.06 76.72 83.97
Table 2: Evaluation of pixel-level segmentation accuracy on extended BTAD data using real-time Segformer, BiseNet V2 and STDC.
Backbone Category CutPaste DRAEM GLASS DFMGAN RealNet AnomalyDiffusion FAST
mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$
Segformer 01 66.94 78.20 67.86 80.14 68.02 79.57 67.02 78.03 67.17 80.20 66.55 76.31 75.93 86.12
02 65.04 83.64 69.52 82.96 69.99 83.58 68.75 84.92 70.64 83.90 68.06 84.74 70.63 81.63
03 50.96 60.41 50.39 54.30 51.77 53.53 38.95 41.55 48.76 57.50 54.85 80.20 79.40 85.64
BiseNet V2 01 57.15 69.88 49.16 63.48 44.09 50.57 49.49 59.20 45.45 57.65 46.66 55.18 58.74 68.98
02 59.45 82.05 66.46 80.29 66.37 79.46 66.02 79.21 66.11 81.67 65.57 84.00 68.02 82.40
03 31.84 40.62 36.15 39.04 30.80 37.15 20.12 21.48 29.55 33.11 42.27 74.41 77.87 92.49
STDC 01 48.06 59.86 42.17 65.36 45.51 60.12 44.68 51.71 32.91 49.21 44.85 55.29 44.95 53.47
02 59.80 77.57 64.96 84.32 65.02 81.94 64.85 75.32 64.00 82.64 64.73 78.93 67.76 82.16
03 19.76 25.20 36.14 38.80 17.04 28.01 14.67 16.55 22.57 24.79 41.71 65.45 84.04 92.36
Anomaly Segmentation Table. 1 and 2 report pixel-level segmentation results on various datasets using Segformer trained with FAST-augmented data. We observe that FAST achieves an average mIoU of 76.72% and accuracy of 83.97%, significantly outperforming the strongest prior method, DRAEM (74.75% Acc), by 9.22 points, respectively. Improvements are particularly notable in challenging categories: in capsule, FAST increases mIoU from 51.39% (DRAEM) to 63.22% (↑11.83); on grid, from 47.75% to 52.45% (↑4.70); and on transistor, from 84.22% to 91.80% (↑7.58). Even in relatively easier categories such as bottle and tile, FAST still yields consistent improvements of 7.35 and 0.92 mIoU points, respectively. These results demonstrate that the combination of mask-aware noise injection via FARM and coarse-to-fine accelerated sampling via AIES enables more realistic and structurally coherent anomaly synthesis, leading to superior segmentation performance. Similar trends are observed when replacing Segformer with other real-time backbones such as BiseNetV2 and STDC, as shown in Supplementary Materials A.6, confirming the generalizability of FAST across different segmentation architectures.
Figure 3: Visualization results of different anomaly synthesis methods on the MVTec dataset. Columns correspond to synthesis methods (from left to right: MVTec AD, CutPaste, DRAEM, GLASS, DFMGAN, RealNet, Anomaly Diffusion, FAST), and rows correspond to product categories (from top to bottom: hazel_nut, transistor, tile, metal_nut, cable).
Qualitative Comparison. Fig. 3 visually compares anomaly samples synthesized by different anomaly synthesis methods across several MVTec-AD categories. It can be observed that traditional unsupervised methods such as CutPaste and DRAEM generate anomalies by overlaying arbitrary textures or patches without any semantic guidance. For instance, in the Cable category, anomalies produced by CutPaste appear as artificial, block-like overlays lacking meaningful texture or structure. Similarly, DRAEM and GLASS introduce unrealistic color distortions and incoherent patterns in the Transistor category, which deviate significantly from typical industrial anomalies. DL-based approaches (DFMGAN, RealNet, and AnomalyDiffusion) generate more visually plausible results, but still exhibit noticeable shortcomings. For instance, RealNet often introduces color shifts and boundary artifacts, as seen in the tile and cable cases, where anomalies appear overly smooth or blurred. DFMGAN and AnomalyDiffusion are able to synthesize more coherent shapes (e.g., spray-paint-like anomalies in hazel_nut), yet they suffer from inaccurate boundaries and structural mismatches, as is especially evident in the tile (AnomalyDiffusion) and cable (DFMGAN) categories. In contrast, FAST consistently produces anomalies that closely resemble realistic defects while maintaining precise alignment with the annotated masks. In the metal_nut and hazel_nut cases, FAST is the only method that preserves both texture and shape fidelity within the intended regions, demonstrating superior controllability and structural consistency. These results validate the effectiveness of the proposed FAST in segmentation-oriented anomaly synthesis.
4.3 Ablation Studies
Figure 4: SIAS results with other sampling strategies. Columns correspond to sampling strategies (from left to right: ground truth, DDPM (1000 steps), DDIM (50 steps), PLMS (50 steps), AIAS (50 steps), and rows correspond to categories (from top to bottom: hazelnut, pill, screw). Further qualitative results (trained on MVTec and BTAD) are provided in the Supplementary Materials A.8.
The Impact of AIAS. We compare our proposed AIAS strategy with several widely-used training-free samplers, including DDPM ho2020denoising with 1000 steps, DDIM song2021denoising with 50 steps and PLMS liu2022pseudo with 50 steps. These methods represent state-of-the-art discrete-time sampling approaches for diffusion-based models. To ensure fairness, we exclude continuous-time solvers, as they rely on a fundamentally different formulation based on ODEs or SDEs, which necessitates a distinct training paradigm and architectural adjustments incompatible with our discrete-time framework. Quantitative results are reported in Fig. 5. While DDPM achieves competitive results on certain categories (e.g., carpet, tile), it requires 1000 iterative steps, making it over 20× slower than AIAS in practice. DDIM and PLMS, though more efficient, exhibit inconsistent performance across categories and often underperform AIAS, particularly on challenging textures such as capsule, grid, and transistor. In contrast, AIAS achieves the best results on the majority of categories and consistently provides competitive or superior performance in both mIoU and accuracy, demonstrating its ability to generate segmentation-aligned anomalies with significantly fewer steps. It further indicates that by analytically aggregating multiple DDPM transitions into coarse-to-fine segments, AIAS reduces the discretization error inherent in single-step samplers (e.g., DDIM) or fixed multistep solvers (e.g., PLMS), allowing a closer approximation of the true posterior within just 50 steps. Fig. 4 further illustrates the qualitative advantage. For example, in the hazel_nut class, the anomalies produced by DDPM, DDIM, and PLMS display noticeable color inconsistencies near the anomaly boundary, resulting from distributional mismatch with the background. In comparison, FAST-produced anomalies that are well blended into the context, with sharper and more realistic structural alignment.
Figure 5: The effect of different sampling methods on SIAS in the MVTec dataset. Top row shows per-category segmentation performance using mIoU; bottom row shows performance using Acc. Detailed per-category results of AIAS are reported in Supplementary Material A.6.
Table 3: Comparison of pixel-level anomaly segmentation using different steps on the MVTec dataset.
Category Step 2 Step 5 Step 10 Step 30 Step 50 Step 100 Step 200 Step 500 Step 1000
mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$
bottle 77.03 80.96 80.55 85.08 83.26 85.90 84.59 87.89 86.86 90.90 83.75 86.95 84.04 88.54 83.52 88.19 81.65 84.83
cable 47.39 48.66 69.58 73.11 71.23 75.07 73.34 77.59 73.71 77.94 72.99 76.50 72.83 76.51 75.23 79.32 73.45 78.06
capsule 43.56 48.58 49.81 54.22 54.85 59.31 61.12 67.08 63.22 71.12 63.15 71.17 62.12 71.76 62.83 70.88 60.01 66.87
carpet 70.24 80.98 73.22 83.18 73.10 84.06 73.56 80.50 73.84 83.53 73.41 82.92 73.17 81.90 73.27 82.49 75.99 84.14
grid 48.15 61.75 50.03 63.28 50.89 71.35 48.76 61.17 52.45 70.70 50.03 65.41 52.06 67.28 49.18 63.63 50.91 63.19
hazel_nut 76.16 78.75 84.16 86.50 90.45 94.04 90.49 94.04 90.81 94.79 90.82 94.16 90.87 94.27 90.77 94.71 89.81 93.31
leather 62.11 66.86 66.74 76.16 67.09 76.51 65.44 72.41 66.60 74.18 66.88 74.22 65.87 87.88 67.95 83.62 71.03 80.32
metal_nut 92.06 93.57 93.94 95.72 94.71 96.98 94.47 96.31 94.65 96.88 94.74 97.19 94.50 96.59 94.72 96.80 94.63 97.18
pill 50.03 55.46 80.01 82.53 90.07 93.80 90.02 94.24 90.17 94.07 89.82 94.10 89.80 93.22 90.15 94.34 89.36 93.79
screw 46.07 52.01 47.92 56.55 50.04 56.21 50.11 60.85 49.94 57.48 50.06 58.66 48.41 61.05 47.71 54.90 49.35 59.18
tile 87.26 93.92 89.46 94.96 89.72 93.92 89.58 93.68 90.13 93.77 89.93 94.45 90.02 93.73 89.71 93.38 91.01 94.72
toothbrush 58.54 67.15 76.65 87.41 76.96 90.29 74.36 90.78 74.98 88.63 74.17 87.29 73.32 86.49 75.66 89.50 76.10 91.25
transistor 66.42 71.59 66.08 70.23 77.27 79.66 89.45 92.65 91.80 94.50 91.39 94.66 89.67 93.50 90.32 93.21 89.59 93.41
wood 68.69 78.28 74.23 81.07 75.97 81.18 78.76 84.99 78.77 86.31 77.00 83.95 77.60 82.85 77.71 83.45 80.03 85.30
zipper 68.85 75.26 70.92 81.44 72.44 84.99 73.08 81.91 72.80 84.73 71.99 81.94 71.71 82.21 71.73 83.47 72.45 82.35
Average 64.17 70.25 71.55 78.10 74.54 81.55 75.81 82.41 76.72 83.97 76.01 82.90 75.73 83.85 76.03 83.46 76.36 83.19
Figure 6: Segmentation-oriented industrial anomaly synthesis results at different steps of AIAS. Columns correspond to increasing sampling steps $T$ (from left to right), and rows correspond to product categories (from top to bottom: capsule, bottle, carpet).
The Impact of AIAS under different step sizes. We further investigate the segmentation performance of AIAS under varying numbers of reverse steps, ranging from 2 to 1000, as reported in Table. 3. Remarkably, AIAS approximates the performance of full-step DDPM using only 10 steps, and reaches near-optimal results by 50 steps, demonstrating the effectiveness of our coarse-to-fine aggregation strategy. Performance improves rapidly as $t$ increases from 2 to 50, since early segments capture the global layout and coarse structure of anomalies, which are most relevant for segmentation. This trend is also visually confirmed in Fig. 6. Beyond this point, performance gains gradually saturate, indicating that additional steps primarily refine high-frequency details with limited impact on segmentation accuracy. Notably, when $t=1000$, AISA degenerates to the original DDPM sampling process, where each segment $t_{e},t_{s}]$ corresponds to a single denoising step. The convergence of performance at this point validates that our multi-step analytical updates provide a faithful approximation of the full diffusion trajectory, preserving both global semantics and fine-grained anomaly cues while significantly reducing sampling cost. Furthermore, excessive denoising steps may introduce over-smoothing or amplify reconstruction inconsistencies, potentially weakening the alignment between synthesized anomalies and segmentation-relevant structures. Overall, these results highlight that AIAS not only accelerates sampling, but also introduces an inductive structural bias beneficial for anomaly segmentation. In practice, the optimal balance between quality and efficiency is achieved within 10–50 steps.
Figure 7: Qualitative ablation results with and without FARM on MVtec dataset. Columns correspond to category–anomaly pairs (from left to right: Wood_color, Tile_oil, Hazel_nut_crack, Cable_missing_wire; and rows correspond to sampling configurations (from top to bottom: without FARM (w/o FARM) and with FARM (w/ FARM).
The Impact of FARM. To evaluate the effectiveness of FARM, we conduct an ablation study by comparing the model’s performance with (w/ FARM) and without (w/o FARM) FARM under identical AIAS settings. Results on the MVTec dataset are reported in Fig. 8. The inclusion of FARM leads to substantial improvements in segmentation performance, with average mIoU increasing from 65.33 to 76.42 and accuracy increasing from 71.24 to 83.97. The performance gains are particularly pronounced in challenging categories characterized by fine-grained or complex structures, such as capsule (↑14.1 mIoU), grid (↑14.7 mIoU), and transistor (↑29.5 mIoU). Even in relatively easier categories like tile and hazel_nut, FARM consistently enhances accuracy and boundary localization, as shown in Fig. 7. More detailed analysis of FARM can be found in Supplementary Material A.7.
Figure 8: Qualitative ablation results with and without FARM on MVtec dataset. Columns correspond to product categories and rows correspond to mIou and Acc). Detailed per-category results for ablation study of FARM are reported in Supplementary Material A.6.
5 Conclusion
In this work, we proposed FAST, a segmentation-oriented foreground-aware diffusion framework tailored for anomaly synthesis. To address the limitations of existing anomaly synthesis methods, specifically their limited controllability and lack of structural awareness, we introduced two key components: the Foreground-Aware Reconstruction Module (FARM), which adaptively injects anomaly-aware noise at each sampling step, and the Anomaly-Informed Efficient Sampling (AIAS), a training-free strategy that accelerates sampling via coarse-to-fine aggregation. Built upon a discrete-time latent diffusion backbone, FAST enables the synthesis of segmentation-aligned anomalies with as few as 10 denoising steps. Extensive experiments on MVTec-AD and BTAD demonstrate that FAST outperforms existing baselines in downstream segmentation. FAST represents a promising step toward controllable and efficient segmentation-oriented industrial anomaly synthesis.
Appendix A Supplementary Materials
A.1 Proof of Lemma 1
Lemma 1 [Linear–Gaussian closure] Let $\{x_{k}\}_{k=0}^{K}\subset\mathbb{R}^{d}$ satisfy the recursion
$x_{k-1}\;=\;C_{k}\,x_{k}\;+\;d_{k}\;+\;\varepsilon_{k},\quad\varepsilon_{k}\sim\mathcal{N}(0,\Sigma_{k}),\quad\varepsilon_{k}\!\perp\!\!\{\!x_{k},\varepsilon_{k+1},\ldots\!\},$ (13)
where $C_{k}\!\in\!\mathbb{R}^{d\times d}$, $d_{k}\!\in\!\mathbb{R}^{d}$, and $\Sigma_{k}\!\in\!\mathbb{R}^{d\times d}$ are deterministic. Then, for every integer $m$ with $1\!\leq\!m\!\leq\!k$, $x_{k-m}$ is again an affine–Gaussian function of $x_{k}$:
$x_{k-m}\;=\;\underbrace{\Bigl(\prod_{i=0}^{m-1}C_{k-i}\Bigr)}_{=:C^{(m)}_{k}}x_{k}\;+\;\underbrace{\sum_{i=0}^{m-1}\Bigl(\prod_{j=1}^{i}C_{k-j}\Bigr)d_{k-i}}_{=:d^{(m)}_{k}}\;+\;\varepsilon^{(m)}_{k}$, (14)
where
$\varepsilon^{(m)}_{k}\sim\mathcal{N}\!\bigl(0,\Sigma^{(m)}_{k}\bigr),\qquad\Sigma^{(m)}_{k}=\sum_{i=0}^{m-1}\Bigl(\prod_{j=1}^{i}C_{k-j}\Bigr)\Sigma_{k-i}\Bigl(\prod_{j=1}^{i}C_{k-j}\Bigr)^{\!\top}.$ (15)
Proof.
: Base case ($m=1$).
  • Eq. 14 with $m=1$ is exactly the recursion Eq. 13.
Induction step.
  • Assume Eq. 14 and .15 hold for $m=r$ with $1\leq r<k$:
    $x_{k-r}=C^{(r)}_{k}x_{k}+d^{(r)}_{k}+\varepsilon^{(r)}_{k},\qquad\varepsilon^{(r)}_{k}\sim\mathcal{N}\!\bigl(0,\Sigma^{(r)}_{k}\bigr),\qquad\varepsilon^{(r)}_{k}\!\perp\!x_{k}.$
  • Apply Eq. 13 once more:
    $x_{k-(r+1)}$ $=C_{k-r}x_{k-r}+d_{k-r}+\varepsilon_{k-r}$ (16)
    $=C_{k-r}\bigl(C^{(r)}_{k}x_{k}+d^{(r)}_{k}+\varepsilon^{(r)}_{k}\bigr)+d_{k-r}+\varepsilon_{k-r}$
    $=\underbrace{C_{k-r}C^{(r)}_{k}}_{C^{(r+1)}_{k}}\,x_{k}+\underbrace{C_{k-r}d^{(r)}_{k}+d_{k-r}}_{d^{(r+1)}_{k}}+\underbrace{C_{k-r}\varepsilon^{(r)}_{k}+\varepsilon_{k-r}}_{\varepsilon^{(r+1)}_{k}}$.
Since $\varepsilon^{(r)}_{k}$ and $\varepsilon_{k-r}$ are independent zero–mean Gaussians, their linear combination $\varepsilon^{(r+1)}_{k}$ remains Gaussian with covariance $\Sigma^{(r+1)}_{k}=C_{k-r}\Sigma^{(r)}_{k}C_{k-r}^{\!\top}+\Sigma_{k-r},$ exactly matching Eq. 15 for $m=r+1$. Hence the statement holds for all $m$ by induction.
Remark 1.
The empty product convention $\prod_{j=1}^{0}C_{k-j}=I_{d}$ is used in Eq. 14.
A.2 Proof of Lemma 2
Lemma 2 [Closed-form reverse from $\boldsymbol{t}_{s}\to t_{e}$}] Fix indices $0\leq t_{e}<t_{s}\leq T$, and let the single-step coefficients $(A_{t},B_{t},\sigma_{t}^{2})$ be defined as in Eq. 13. Then the aggregated reverse kernel over $t_{s}\to\cdots\to t_{e}$ is affine–Gaussian:
$x_{t_{e}}=\Pi_{t_{e}}^{t_{s}}\,x_{t_{s}}+\Sigma_{t_{e}}^{t_{s}}\,\hat{x}_{0}+\varepsilon_{t_{e}},$ (17)
where
$\Pi_{t_{e}}^{t_{s}}:=\prod_{i=t_{e}+1}^{t_{s}}B_{i},\quad\Sigma_{t_{e}}^{t_{s}}:=\sum_{i=t_{e}+1}^{t_{s}}A_{i}\prod_{j=i+1}^{t_{s}}B_{j},\quad\varepsilon_{t_{e}}\sim\mathcal{N}\!\left(0,\sum_{i=t_{e}+1}^{t_{s}}\left(\prod_{j=i+1}^{t_{s}}B_{j}\right)^{2}\sigma_{i}^{2}\mathbf{I}\right).$
Proof.
Apply Lemma 1 with $C_{k}=B_{k}$, $d_{k}=A_{k}\hat{x}_{0}$, $\Sigma_{k}=\sigma_{k}^{2}\mathbf{I}$, and $m=t_{s}\!-\!t_{e}$. Equations Eq. 17 coincide with the general expressions Eq. 14–Eq. 15, so the result follows directly. ∎
A.3 Loss function
The training objective of FAST consists of two components: the standard denoising loss and the reconstruction loss. The denoising loss encourages accurate noise prediction across all spatial regions, while the reconstruction loss ensures that FARM accurately reconstructs anomaly-only content, and allowes the inserted noise to remain compatible with the global sampling dynamics, thereby preserving the stability of the overall generation process.
$\mathcal{L}_{\text{FAST}}$ $=\lambda_{\text{1}}\cdot\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon},t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right\|_{2}^{2}\right]$ (18)
$\quad+\lambda_{\text{2}}\cdot\mathbb{E}_{\mathbf{x}^{\mathrm{an}}_{0},\mathbf{x}_{t},\mathcal{M}}\left[\left\|F_{\phi}(\mathbf{x}_{t},\mathcal{M},t)-\mathbf{x}^{\mathrm{an}}_{0}\right\|_{2}^{2}\right],$
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ is the reference noise for the denoising target, the $\mathbf{x}^{\mathrm{an}}_{0}$ is the anomaly-only content with pure background, $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ and $F_{\phi}$ denote LDM and FARM, respectively. The scalar weights $\lambda_{\text{1}}$ and $\lambda_{\text{2}}$ balance the contributions of the two losses
A.4 Pseudo-code of AIAS
Algorithm 3 Anomaly-Informed Accelerated Sampling
Input: Mask $\mathcal{M}$, clean background $\mathbf{x}_{\mathrm{full}}^{\mathrm{bg}}$, clean background latent $\mathbf{x}_{0}^{\mathrm{bg}}$, prediction $\hat{\mathbf{x}}_{0}$ from $\epsilon_{\theta}$
boundary schedule $\mathcal{B}=\{t_{1}<t_{2}<\dots<t_{K}=T\}$ and $t_{1}=2$ in our experiments
Output: Syntheised image ${\mathbf{x}}_{\mathrm{full}}$
Initialize noisy latent $\mathbf{x}_{t_{K}}\sim\mathcal{N}(0,\mathbf{I})$
for $k=K$ to 11 do
  $t_{s}\leftarrow t_{k}$, $t_{e}\leftarrow t_{k-1}$
  # Coarse multi-step reverse from $t_{s}\to t_{e}$
  Define cofficients $A_{t}=\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}$, $B_{t}=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}$, and $\sigma_{t}^{2}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$
  Compute $\mu\leftarrow(\prod_{i=t_{e}+1}^{t_{s}}B_{i})\mathbf{x}_{t_{s}}+(\sum_{i=t_{e}+1}^{t_{s}}A_{i}\prod_{j=i+1}^{t_{s}}\!B_{j})\hat{\mathbf{x}}_{0}$
  Sample noise $\varepsilon\sim\mathcal{N}(0,(\sum_{i=t_{e}+1}^{t_{s}}\Bigl(\prod_{j=i+1}^{t_{s}}B_{j}\Bigr)^{\!\!2}\sigma_{i}^{2})\mathbf{I})$
  $\mathbf{x}_{t_{e}}\leftarrow\mu+\varepsilon$
  # Forward diffuse background to $t_{e}$
  $\mathbf{x}_{t_{e}}^{\mathrm{bg}}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t_{e}}}\mathbf{x}_{0}^{\mathrm{bg}},(1-\bar{\alpha}_{t_{e}})\mathbf{I})$
  $\mathbf{x}_{t_{e}}^{R}\leftarrow\mathrm{FARM}(\mathbf{x}_{t_{e}})$
  $\mathbf{x}_{t_{e}}\leftarrow\mathcal{M}\odot\mathbf{x}_{t_{e}}^{R}+(1-\mathcal{M})\odot\mathbf{x}_{t_{e}}^{\mathrm{bg}}$
end for
# Fine posterior refinement
for $t=t_{1}$ down to 0 do
  Predict $\hat{\mathbf{x}}_{0}\leftarrow f_{\theta}(\mathbf{x}_{t},t)$
  ${\mathbf{x}}_{t-1}\leftarrow q(\mathbf{x}_{t-1}\mid\mathbf{x}_{t},\hat{\mathbf{x}}_{0})$
end for
${\mathbf{x}}_{\mathrm{full}}\leftarrow\mathcal{M}\odot Decode(\mathbf{x}_{0})+(1-\mathcal{M})\odot\mathbf{x}_{\mathrm{full}}^{\mathrm{bg}}$
A.5 Training Configuration
To synthesize abnormal data, we utilize the complete set of normal images, their corresponding masks, and associated textual descriptions for each type of anomaly within every category of products. Notably, the original GLASS framework comprises three branches,a normal-sample branch, a feature-level anomaly synthesis branch guided by gradient ascent, and an image-level branch that overlays external textures. Therefore, its output is unsuitable directly for pixel-level anomaly segmentation and other downstream sgementation models. Accordingly, we revised its synthesis process to align with our segmentation-based evaluation protocol. We release the modified implementation together with the FAST to ensure fairness.
  • Model Settings. We set the total number of diffusion steps during training to $T=1000$. For sampling, the range from step 2 to 1000 is uniformly divided into 50 steps, followed by a fine-grained adjustment phase over the initial steps $[0,2]$ to enhance reconstruction fidelity. The model is trained with a batch size of 4 and a learning rate of 1.5e-4. The text embedding $E$ consists of 8 tokens.
  • Prompt Construction. For the MVTec dataset, prompts are formed by appending the anomaly type to the product category name. For BTAD, due to anonymized category labels, we use a generic prompt: “damaged". Textual embeddings follow the protocol of AnomalyDiffusion, where each prompt is tokenized into 8 discrete units and embedded using a pre-trained BERT encoder [4].
  • Hardware and Runtime. All models are trained on a setup of eight NVIDIA A100 GPUs (40GB each), with training proceeding for roughly 80k iterations.
A.6 Other quantitative experiments
We provide extended evaluation results to complement the findings reported in the main manuscript. We present detailed, category-wise performance metrics on the MVTec and BTAD benchmarks, employing BiseNet V2 and STDC as the segmentation backbones. Moreover, we further analyze the influence of different sampling strategies—except our AIAS method—on downstream segmentation performance using Segformer.
All experiments are conducted under identical settings to those used in the main study. The results consistently demonstrate that our proposed FAST framework significantly outperforms existing anomaly synthesis techniques in enhancing segmentation accuracy across diverse categories.
Table 4: Evaluation of pixel-level segmentation accuracy on extended MVTec data using real-time BiseNet V2.
Category CutPaste DRAEM GLASS DFMGAN RealNet AnomalyDiffusion FAST
mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$
bottle 71.77 78.57 75.13 79.17 57.81 60.79 64.28 71.31 72.16 75.55 75.28 85.11 78.48 83.18
cable 46.00 57.08 53.88 60.96 16.63 16.65 57.09 63.25 51.22 62.32 60.55 74.96 70.91 75.77
capsule 25.97 37.04 36.82 42.19 19.53 51.89 28.40 31.18 35.97 39.39 26.77 32.87 48.56 54.22
carpet 58.98 72.22 68.42 77.21 64.77 73.93 62.13 67.98 8.98 9.01 58.18 64.69 68.94 77.20
grid 24.68 44.17 42.81 63.34 6.50 6.91 10.17 15.23 10.61 11.47 18.98 24.30 39.15 51.78
hazel_nut 47.93 53.57 74.83 81.35 71.54 75.62 79.78 84.37 60.16 65.93 57.26 70.41 88.08 93.45
leather 31.11 58.36 55.07 61.58 57.98 71.84 31.77 34.82 53.77 63.85 50.02 61.60 67.18 74.23
metal_nut 82.95 87.73 91.58 94.73 83.82 85.42 91.17 93.57 88.38 90.73 85.52 90.20 93.62 95.82
pill 55.62 67.04 45.23 48.99 23.88 24.15 82.40 84.30 72.59 86.32 80.87 87.02 85.12 89.60
screw 4.88 6.63 25.08 35.77 12.32 13.11 38.14 40.36 22.35 23.78 23.23 29.91 33.49 41.12
tile 76.25 85.75 86.17 90.45 77.32 80.28 85.69 90.12 77.16 84.84 79.32 85.63 86.86 92.12
toothbrush 35.69 50.45 57.66 79.15 38.86 51.97 48.83 58.76 32.38 37.88 44.33 69.32 73.04 87.34
transistor 44.48 51.79 59.88 65.96 44.93 53.04 76.52 82.13 61.68 68.59 76.34 89.94 91.10 93.81
wood 35.51 46.00 49.82 62.09 36.41 51.10 51.84 63.70 47.29 61.35 52.06 72.75 68.15 72.69
zipper 51.61 63.09 66.88 75.75 61.99 70.07 60.61 71.11 66.09 77.54 57.86 67.64 66.59 78.16
Average 46.23 57.30 59.28 67.91 44.95 52.45 57.92 63.48 50.72 57.24 56.44 67.09 70.62 77.37
Table 5: Evaluation of pixel-level segmentation accuracy on extended MVTec data using real-time STDC.
Category CutPaste DRAEM GLASS DFMGAN RealNet AnomalyDiffusion FAST
mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$
bottle 71.37 82.19 73.31 78.23 63.22 69.25 67.66 76.52 69.44 75.68 72.66 84.94 76.82 80.65
cable 42.88 54.74 50.02 58.38 49.38 57.80 57.74 62.86 35.97 38.81 59.43 74.22 54.85 60.26
capsule 21.73 30.72 36.31 41.68 22.91 27.18 25.60 27.96 31.08 34.25 22.90 26.06 49.35 55.29
carpet 50.79 66.68 66.28 76.70 63.18 77.85 58.58 71.83 57.48 68.51 56.16 68.47 64.52 75.02
grid 15.24 25.75 30.29 41.50 19.89 24.72 1.39 1.39 5.37 5.85 16.20 24.63 20.82 25.60
hazel_nut 58.48 65.59 78.75 83.66 68.57 85.83 81.77 84.66 70.16 82.40 61.83 92.42 87.96 93.99
leather 38.12 58.63 44.63 56.84 57.53 73.90 21.29 22.28 36.76 53.88 46.98 59.89 60.38 75.90
metal_nut 81.13 86.63 91.12 94.08 83.97 89.37 90.68 92.73 86.85 91.45 85.81 90.06 93.01 95.32
pill 50.00 60.28 55.47 61.05 44.48 48.11 80.41 82.55 63.96 65.96 78.23 84.35 82.15 86.48
screw 2.80 4.98 16.16 23.05 16.81 19.33 34.93 38.76 17.93 18.76 1.27 2.00 17.82 21.25
tile 69.86 78.18 84.75 91.31 79.86 88.65 85.36 89.72 70.29 77.70 76.96 84.07 86.29 93.89
toothbrush 41.19 52.81 53.72 76.55 37.46 40.91 36.78 38.94 33.85 43.03 35.39 48.93 75.76 87.32
transistor 58.24 68.80 65.57 80.31 62.64 69.32 78.38 87.23 62.57 72.45 71.96 83.28 93.01 96.05
wood 31.75 43.27 55.25 60.82 36.31 45.67 26.36 33.13 37.23 43.37 48.90 62.57 72.27 78.06
zipper 47.51 59.24 61.03 68.53 59.07 69.39 44.42 51.83 60.04 71.52 56.77 66.66 52.03 67.69
Average 45.41 55.90 57.51 66.18 51.02 59.15 52.76 58.81 49.27 56.24 52.76 63.50 65.80 72.85
Table 6: Ablation study of FARM on the MVTec dataset using the real-time Segformer.
Category mIoU (w/o FARM) $\uparrow$ Acc (w/o FARM) $\uparrow$ mIoU (w/ FARM) $\uparrow$ Acc (w/ FARM) $\uparrow$
bottle 80.65 83.46 86.86 90.90
cable 65.99 70.50 73.71 77.94
capsule 49.08 53.25 63.22 71.12
carpet 72.46 80.84 73.84 83.53
grid 37.79 42.61 52.45 70.70
hazelnut 69.20 72.55 90.81 94.79
leather 61.42 65.91 66.60 74.18
metal_nut 89.59 94.31 94.65 96.88
pill 46.73 48.44 90.17 94.07
screw 46.48 54.42 49.94 57.48
tile 88.91 93.28 90.13 93.77
toothbrush 66.29 81.40 74.98 88.63
transistor 62.35 67.46 91.80 94.50
wood 71.55 79.47 78.77 86.31
zipper 71.40 80.76 72.80 84.73
Average 65.33 71.24 76.72 83.97
Table 7: Ablation Study of AIAS with other training-free sampling Methods on MVTec-AD data via Segformer.
Category DDPM (1000 steps) DDIM (50 steps) PLMS (50 steps) AIAS (50 steps)
mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$ mIoU $\uparrow$ Acc $\uparrow$
bottle 81.65 84.83 82.87 86.03 81.49 84.44 86.86 90.90
cable 73.45 78.06 74.21 78.41 74.78 78.91 73.71 77.94
capsule 60.01 66.87 58.02 64.03 56.92 61.90 63.22 71.12
carpet 75.99 84.14 75.33 83.58 75.41 82.39 73.84 83.53
grid 50.91 63.19 50.85 67.91 50.43 61.42 52.45 70.70
hazel_nut 89.81 93.31 89.69 93.03 89.42 92.96 90.81 94.79
leather 71.03 80.32 66.00 72.48 71.85 81.47 66.60 74.18
metal_nut 94.63 97.18 94.50 96.47 93.93 96.69 94.65 96.88
pill 89.36 93.79 89.84 93.03 89.93 93.66 90.17 94.07
screw 49.35 59.18 48.89 57.26 48.78 55.62 49.94 57.48
tile 91.01 94.72 89.23 92.90 89.96 93.25 90.13 93.77
toothbrush 76.10 91.25 74.79 88.48 76.02 91.00 74.98 88.63
transistor 89.59 93.41 89.35 92.37 89.17 91.99 91.80 94.50
wood 80.03 85.30 79.29 84.03 79.61 84.65 78.77 86.31
zipper 72.45 82.35 71.01 83.00 72.06 81.02 72.80 84.73
Average 76.36 83.19 75.59 82.20 75.98 82.09 76.72 83.97
A.7 More analysis of FARM
These improvements of FARM are not only empirically significant, but also consistent with intuitive understanding. Without FARM, the segmentation-oriented industrial anomaly synthesis relies solely on frozen pre-trained weights and weak conditioning from learned textual embeddings. This limits the model’s ability to capture the structural characteristics of industrial anomalies, often leading to visually perturbed but semantically uninformative results. In contrast, FARM explicitly reconstructs anomaly-only content from noisy latents and produce spatially localized, anomaly-aware noise into the sampling process. Additionally, by incorporating both spatial masking and timestep encoding, FARM guides the model to focus on abnormal regions—information that would otherwise be uniformly treated in the absence of FARM. Together, these mechanisms improve the structural fidelity, localization precision, and segmentation relevance of synthesized anomalies.
A.8 Other qualitative experiments
We also provide additional qualitative results to supplement the main paper. Specifically, we present synthesized anomalies across multiple categories from MVTec and BTAD, along with comparisons against CutPaste, DRAEM, GLASS, RealNet, DFMGAN, and AnomalyDiffusion. Each figure includes both the generated images and their corresponding segmentation masks.
Figure 9: Visualization results of different anomaly synthesis methods on the MVTec dataset. Columns correspond to synthesis methods (from left to right: MVTec AD, CutPaste, DRAEM, GLASS, DFMGAN, RealNet, Anomaly Diffusion, FAST), and rows correspond to product categories (from top to bottom: bottle, cable, capsule, carpet, grid).
Figure 10: Visualization results of different anomaly synthesis methods on the MVTec dataset. Columns correspond to synthesis methods (from left to right: MVTec AD, CutPaste, DRAEM, GLASS, DFMGAN, RealNet, Anomaly Diffusion, FAST), and rows correspond to product categories (from top to bottom: hazel_nut, leather, metal_nut, pill, screw).
Figure 11: Visualization results of different anomaly synthesis methods on the BTAD dataset. Columns correspond to synthesis methods (from left to right: MVTec AD, CutPaste, DRAEM, GLASS, DFMGAN, RealNet, Anomaly Diffusion, FAST), and rows correspond to product categories.